OpenStudy (anonymous):

During 2003, a share of stock in the Coca-Cola company sold for $39. During 2008. the price hit $56 per share. Express the increase in price as a percent of the price in 2003. Round to the nearest tenth

OpenStudy (compassionate):

Will I get a medal for answering this?

OpenStudy (anonymous):

well duh if you help me :)

OpenStudy (anonymous):

Rate increase = (54-39)/(2008-2003)

OpenStudy (compassionate):

One second.

OpenStudy (anonymous):

Sorry its 56

OpenStudy (compassionate):

You sniped me :(

OpenStudy (anonymous):

38.7 percent?

OpenStudy (anonymous):

Rate increase=(56-39)/(2008-2003) = 3.4 Now, percentage increase = 3.4 * 100% =34%

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

oh okay thank you! :)

OpenStudy (anonymous):

Welcome